OpenStudy (anonymous):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (anonymous):

@Nnesha

Directrix (directrix):

To help get started, the set {–1, 1} is not closed under addition. For a set to closed under an operation means that, in this case, if you add any two elements of the set, then you have to get an element of the set as the sum.

Directrix (directrix):

Choose 1 and -1 from the set. Add them to get zero. Zero is not in the set. Therefore, the set is not closed under additon.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

B and D?

Directrix (directrix):

Now, substraction. If you subtract any two members of the set, will you always get back a member of the set? You said B so you're thinking yet. Let's see.

OpenStudy (anonymous):

kk

Directrix (directrix):

I am taking 1 from the set and subtracting it from itself. 1-1 =0 which is not in the set. Not closed under subtraction.

Directrix (directrix):

What about this: multiplication . Closed? What do you think?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Is b and c right

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

:(

Directrix (directrix):

Did you change your mind about the set being closed under multiplication?

OpenStudy (anonymous):

Ya

Directrix (directrix):

-1 times 1 = -1 1 times 1 = 1 -1 times -1 = 1 Closed under multiplication

Directrix (directrix):

That leaves division.

Directrix (directrix):

What do you think, closed or not under division?

OpenStudy (anonymous):

I dont think it is

Directrix (directrix):

Can you give an example of one number of the set divided by another which would give a number not in the set?

Directrix (directrix):

1/1=1 1/(-1)=-1

Directrix (directrix):

The set appears to be closed under both multiplication and division.

OpenStudy (anonymous):

ohhh i get it!

OpenStudy (anonymous):

Could You help with another?

Directrix (directrix):

Alright, that is good. In making your closed or not closed decision, test some numbers in the set first. Okay?

Directrix (directrix):

>>Could You help with another? Yes. Please post that question in a new thread because this one is so long already.

OpenStudy (anonymous):

ook